# Decentralized Reliability Estimation for Mixnets

### v1

@@ 4.1.2 Adversarial setting
We now evaluate the proposed node performance estimation protocols in adversarial conditions. We consider that the adversary controls a set $A$ of malicious nodes and wants to attack a set $T$ of honest reliable target nodes, where $|A|$ and $|T|$ range from one node to up to $40\%$ of two mixnet layers ($32$ out of $80$ nodes in each layer). The details of the experimental setup can be found in Appendix D.6.
@@
To attack a target node, an adversary node needs to share an edge with the target, either as predecessor or as successor. If the adversary is the successor of the target, then it simply drops packets coming from a target without registering them in its Bloom filter. If the adversary is the predecessor of a target, then it registers the received packet tags in the Bloom filter but then drops packets directed to the target, instead of actually sending them. The purpose of the attack is to degrade the measured performance of target nodes compared to their true performance. Thus, the adversary is more successful as the protocol more heavily underestimates the true performance of target nodes. The cost of the attack for the adversary is an ‘opportunity cost’ given by measured performance degradation suffered by adversarial nodes due to dropping packets for the attack. The attack is thus more costly as the protocol estimates a lower performance for adversarial nodes, compared to the performance that those nodes would have if they did not engage in an attack.
@@
We run over five hundred simulations with various combinations of $|A|$ and $|T|$. In each experiment we compute the adversarial cost for the $|A|$ malicious nodes $c_{A}=|A|-\sum_{i\in A}\hat{\rho}_{i}$ and the cost imposed on the $|T|$ targets $c_{T}=|T|-\sum_{j\in T}\hat{\rho}_{j}$. Note that the overall attack scale increases proportionally to $|A|$x$|T|$, as this determines the number of links where packets are dropped and consequently the magnitude of the attack costs $c_{A}$ and $c_{T}$. This results in $c_{A}$ and $c_{T}$ values ranging from below $10^{-2}$ to above $10$. We use log scale to represent results for the full range of values from small to large.
@@
Figure 3: Results for adversarial settings where all honest nodes are reliable. Each sample represents the aggregate performance penalties $c_{T}$ for $1\leq|T|\leq 64$ targets ($y$ axis) and $c_{A}$ for $1\leq|A|\leq 64$ adversaries ($x$ axis) in each simulation run (total $510$ runs).
@@
In Figure 3 we show the result of each simulation as a dot with coordinates $(c_{A},c_{T})$, As we can see, all the results are close to the diagonal $c_{A}=c_{T}$, meaning that by engaging in this attack a set of malicious nodes take a combined performance penalty similar to the one they impose on the set of targets, even in scenarios where $40\%$ of the targets’ predecessors and/or successors are adversarial. In the experiments shown in this section we consider that all honest nodes have perfect reliability and that all the packet drops are due to the attack. Appendix D.7 shows additional results where the network includes both adversarial and honest but unreliable nodes.
@@ 4.2 Protocol overhead evaluation
In this section we evaluate the overhead introduced by the proposed protocols in terms of the amount of data that needs to be broadcast in each epoch, the data that needs to be stored long-term in a blockchain, and the computational overhead of VRF-based routing. We consider that a number of variables necessary for our protocols are already publicly available for the basic functioning of the mixnet, even before any performance measurement protocols are in place. This includes: parameters defining the mixnet topology, per-epoch selection and assignment of mix nodes to layers, node public keys (both for signing broadcast information and for decryption of packets), node staking and rewards, client subscriptions and information on spent client credentials. We thus focus on the additional overhead introduced by the proposed performance estimation protocol, summarized in Table 2.
We consider a setup with $100$ million packets routed per epoch, of which $1$ million ($1\%$) are measurement packets. The mixnet includes $80$ gateways and $240$ mix nodes divided in $3$ layers of $80$ nodes. The data that needs to be broadcast in each epoch for computing performance scores includes:
@@
We consider that the computed performance scores for all nodes and gateways are stored long-term in the blockchain to act as a historical record of their measured performance. Considering a precision of four decimal places, this requires $2$ bytes per node and gateway, resulting in $640$ bytes for $240$ nodes plus $80$ gateways each epoch. Additional measurement protocol parameters that need to be available in the blockchain, such as $p_{\mathbf{lot}}$ and $\bar{\tau}$, may take a few extra bytes but do not need to be updated every epoch.
@@
Node performance scores 2 B 320 640 B
Table 2: Overhead performance measurement protocol
@@ 5 Related Work
Mixnets for voting applications. A key use case for mixnets in the literature is voting applications [27, 17, 35, 20], where the mixnet must ensure both the unlinkability of voters to their votes, as well as the integrity of the final tally. Therefore, in addition to privacy properties (input-output unlinkability), these solutions offer very high levels of assurance in terms of verifiability of correct behaviour by all participants, e.g.: each client is authorized by a credential to send a single well-formed ballot and verified to do so, while intermediary mix nodes are guaranteed to correctly process every packet without dropping, adding, or substituting any packets. This level of assurance is however achieved at a significant cost per routed packet: first, all transmissions between any parties happen “in public” via a broadcast channel (or “bulletin board”), so that everyone can verify who sent which (encrypted) packets at each step and no one can deny having received any packets (because they are all public); furthermore, every mix node must also produce and publish NIZK-based shuffle proofs of correctness for its published output packets [21], so that those outputs are considered valid and further processed by the next mix node along the route. The end-to-end latency and cost of this approach are acceptable in voting applications but is impractical for general-purpose internet access and other applications,
Mixnets for micro-blogging. Atom [22] is a mixnet designed to send short messages (as in micro-blogging) that provides similar reliability properties using lightweight traps. Atom can route one million $32-$byte messages in $28$ minutes, with latency growing with the number of routed messages and with message size. This is in contrast to the setup considered in this paper, where we expect the mixnet to route client messages with a size in (tens or hundreds of) KB within a second, independently of the number of client messages being routed. This is achieved in our design because the proposed performance estimation protocols are executed a posteriori, after all the messages of an epoch have been delivered, and thus have no impact on the processing time needed to deliver packets compared to the same mixnet not running our performance estimation protocols. Although Atom’s amount of broadcast data is not explicitly provided, given its design, the amount of required broadcast data grows with the number of client messages, in contrast to our approach, where the broadcast overhead remains constant for a given network size and measurement accuracy.
@@
Mixnet reliability. Finally, we review prior work that specifically tackles reliability in mixnets. Early research includes the proposal by Danezis and Sassaman [8] of having mix nodes send heartbeat traffic to themselves through the mixnet, similar to the loops of cover traffic implemented in Loopix, to detect active attacks by malicious nodes that drop or substitute messages. The idea of using reputation to increase the reliability of mixnets was already put forward in early work by Dingledine et al. [11], who proposed a solution that relies on semi-trusted third-party witnesses to verify whether mix nodes are correctly processing messages and help clients avoid under-performing nodes when choosing packet paths. This work however explicitly excludes active adversaries that try to compromise reliability measurements. Follow-up work by Dingledine and Syverson [13] eliminates the need for trusted witnesses by using test measurements and failure reports linked to reputation, as well as commitments to communal randomness, to build mix cascades that can detect dropped or substituted packets. This solution however suffers from “creeping death” attacks on measurements enabled by adversaries biasing the measurement protocol in their favor, such that over time honest nodes lose more reputation than adversarial nodes. Note that our solution avoids this problem by penalizing adversarial nodes symmetrically with respect to the reputational damage inflicted on their targets. Finally, more recent proposals by Leibowitz et al. [25] rely on honest mix nodes having perfect reliability (any packet loss is considered malicious), which is unrealistic in practical scenarios where honest participants may sometimes experience downtime or congestion. The goal of this solution is also slightly different from ours: to disconnect nodes that drop packets rather than to estimate their reliability in a decentralized manner.
6 Conclusion and Future Research
@@
We have introduced a system where reliability estimates can be computed in an efficient and decentralized manner for a continuous mixnet’s links and nodes, given client credentials that authorize sending packets through the mixnet. To realize this concept, we put forward a link performance estimation protocol that uses covert measurement packets to estimate link performances for legitimate packets across the mixnet, and the associated primitive of VRF-based routing that ensures that legitimate client packets have routes chosen according to the mixnet’s routing policy, with malicious clients being unable to bias their packets. We further demonstrate how to use the link estimates as input to a node performance estimation protocol that allows mixnet participants to identify the under-performers amongst them.
@@
Besides formally arguing the security properties of our new primitive, we have empirically evaluated via simulations the practical accuracy of the proposed protocols. To make our proposal practically relevant and easy to integrate with the currently deployed Nym mixnet, we consider $3$ layers and $80$ mix nodes per layer. We find that for this network size one million measurements are sufficient to provide accurate estimations for performance, with the number of required measurements remaining constant regardless of how many legitimate packets are routed. Our novel VRF-based routing primitive is however applicable to any other decryption mixnet with the two main required functionalities, namely issuance of credentials for authorized usage and a broadcast channel, e.g., a blockchain, available to all participants.
@@
In comparison to most verifiable solutions, our solution is compatible with low-latency continuous time mix networks such as Nym as the verification steps happen a posteriori and thus do not add latency to the delivery of packets. Furthermore, the broadcast channel is used for a small fraction of packets ($1\%$ or less if the network routes several hundred million packets or more) rather than all of them. In addition, for each measurement packet only a few parameter values need to be broadcast to enable verifying the packet header, rather than requiring (as in e-voting oriented solutions) storing the entire packet including its payload, recorded repeatedly at each step of its route.
Our work can be extended in various ways. Future research includes protocols for free riding detection and additional mechanisms to protect against fully malicious gateways (rather than utility maximizer adversaries) that e.g., allow their clients to send (illegitimate) traffic without having a valid credential. Considering that the ratio of total to measurement packets received by each node provides information on whether a predecessor has only forwarded legitimate traffic, measurement packets can be used not only to estimate link and node performance, but also to detect traffic injection by mix nodes and gateways and uncover participants that engage in free riding or denial-of-service attacks. Blocklisting policies (and consequent loss of rewards and reputation) for mix nodes and gateways caught injecting traffic can be put in place to discourage malicious behaviour, though we note that these mechanisms may be social as well as cryptographic in nature. Finally, the use of digital signatures by gateways on client packets that act as receipt for clients can facilitate the public detection of malicious gateways that drop clients’ packets (even at the cost of their own utility, i.e., profit), by providing clients with evidence (in the form of a “promise to forward”) that can be verified by others. Note that clients can already use dummy loops [29] in our design to detect Byzantine dropping behavior by gateways, as the Bloom filter mechanism used in our scheme allows clients to trace their own packets (for which they know all packet tags) while maintaining unlinkability for all other clients.
Use of AI-based tools:
This work is entirely human-written without any assistance from AI-based tools.
Appendix A Sphinx packets
Sphinx packets [7] are of the form $\alpha,\beta,\gamma,\delta$ where $\langle\alpha,\beta,\gamma\rangle$ is the header and $\delta$ is the packet’s payload.
  • $\alpha$ is a group element $g^{x}$. It is used to derive a master packet key from the Diffie Hellman operation on $\alpha$ and the recipient’s public-key.
  • $\beta$ is a stream cipher ciphertext encrypted by extending a key derived by the master packet key. It hides the $\beta^{\prime},\gamma^{\prime}$ part of the header for the next hop as well as the identity of the next recipient.
  • $\gamma$ is a MAC signing $\beta$ with a key derived from the master packet key. At each processing step, the recipient verifier that the $\gamma$ is a correct MAC on $\beta$.
  • $\delta$ is a ciphertext encrypting the payload to be forwarded to the next hop. A symmetric cipher is used with a key derived from the master packet key.
When constructing a Sphinx packet one needs to determine the sequence of nodes $n_{1},\ldots,n_{\nu}$ for the path of the packet. Then the headers $\langle\alpha_{i},\beta_{i},\gamma_{i}\rangle$ are calculated. One key operation is that $\alpha_{1}=g^{x}$, while for $i>1$, $\alpha_{i}=(\alpha_{i-1})^{b_{i-1}}$ with $b_{i}$ a blinding value calculated based on the master key of the $i$-th hop. It follows that the $\alpha$ values of the headers are calculated for $i=1,\ldots,\nu$. Once they are determined, the values $\beta_{i},\gamma_{i},\delta_{i}$ are calculated in the reverse order for $i=\nu,\ldots,1$.
Appendix B Proof of Theorem 1
@@
We show how to incorporate a challenge for the Decisional Diffie Hellman assumption $\langle g,a,b,c\rangle$ into the encoding calculation. We set the credential key $\alpha$ to be $a$ and the node public-key $y_{i}$ to be equal to $b$. It follows that the value $\tilde{r}_{i}$ can be calculated as a function of the DDH challenge by $H(\mathsf{rnd},c^{e})$ where $e=b_{0}\ldots b_{i-1}\cdot r_{\mathrm{pkt}}$. Conditional on $e\neq 0$, it follows that $\tilde{r}_{i}$ is uniformly random, unless the adversary queries the value $c^{e}$ to the random oracle. Note that this event can only happen with non-negligible probability in the case the challenge $\langle g,a,b,c\rangle$ is a DDH tuple (as in the other case, the value $c$ is random and hence can only be predicted with negligible probability by any polynomial-time algorithm). As a result, if the adversary queries $c^{e}$ we can readily build a distinguisher by examining the random oracle queries of the adversary. Specifically, the distinguisher can test each query $q$ for $q=c^{e}$ and output $1$ if this is the case while producing 00 otherwise. It is easy to see that this algorithm will produce $1$ with non-negligible probability when the tuple $\langle g,a,b,c\rangle$ follows the DDH distribution while it will produce 00 with overwhelming probability whenever $\langle g,a,b,c\rangle$ is a random tuple. Finally, observe that the probability that $e=0$ is negligible given the randomness of the constituent values in the random oracle model.
The above argument generalizes via a hybrid argument in a straightforward manner to the case that there are multiple clients sending multiple packets. Note that it is in this step that the VRF property is needed since we want to ensure that the $\alpha_{0}$ values of different packets appear independently sampled from the base group. The pseudorandomness property of the underlying function provides us exactly that: the value $r_{\mathsf{pkt}}$ would randomize $\alpha$ within the base group enabling us to repeat the DDH distinguishing argument for each packet.
@@
(II) The statement asks for the pseudorandomness of the values $r_{0},\ldots,r_{\nu-1}$, even in the case that the values $vk,\alpha,y_{0}$ are adversarially chosen. Note that in this case we cannot rely on the DDH assumption or the properties of the VRF. Instead the key observation is that each value $r_{i}$ depends on the public nonce value via the random oracle as well as some other value $s_{i}$ that as long as it is unique, $r_{i}$ would be an independently sampled value. To prove this we consider the event that there are two values $s=g^{e},s^{\prime}=g^{e^{\prime}}$ in this sequence that are equal, i.e., $e=e^{\prime}$. Depending on the step when these values have been calculated the exponents will be a product of a number of values some of which are committed by the adversary, e.g., the discrete-log of the key of an adversarial gateway, the discrete-log of a credential’s randomization key $\alpha$ as well as values produced by random oracle evaluations of such values e.g., the blinding exponents $b_{0},b_{1},b_{2},\ldots$ and the keys of the mix-nodes. Conditional on no collisions in the random oracle $H(\cdot)$ or the VRF function, all these individual values involved in $e,e^{\prime}$ are distinct and, in case they are not outputs of the random oracle, are committed prior to the production of $\mathsf{nonce}$. Based on this it follows that the event $e=e^{\prime}$ would imply the event that $z_{1}\cdot z_{k}\cdot h_{1}\ldots h_{t}=1$ where $h_{1},\ldots,h_{t}$ are random oracle outputs that depend on $\mathsf{nonce}$ and $z_{1},\ldots,z_{k}$ values selected by the adversary. It is easy to see via generalized birthday problem [34] that this would require subexponential overhead in the worst-case (note in our setting $t$ is a small constant).
(III) This statement follows easily due to the fact that the only difference between measurement and non-measurement packets is in the calculation of the $\alpha_{0}$ value: in the case of a non-measurement packet this value is equal to $\alpha^{r_{\mathsf{pkt}}}$, while in the case of a measurement packet it is equal to $g^{r_{\mathsf{pkt}}}$. Given that client and gateway is assumed honest in this case, the exponentiation by the (hidden) value $r_{\mathsf{pkt}}$ results in a full randomization of the resulting element within the base group (with the precondition $\alpha\neq 1$); as a result the two cases are indistinguishable assuming the security of the VRF function. ∎
Appendix C Approaches to setting $\hat{\beta}_{e}$

### v2

@@ 4.1.2 Adversarial setting
We now evaluate the proposed node reliability estimation protocols in adversarial conditions. We consider that the adversary controls a set $A$ of malicious nodes and wants to attack a set $T$ of honest reliable target nodes, where $|A|$ and $|T|$ range from one node to up to $40\%$ of two mixnet layers ($32$ out of $80$ nodes in each layer). The details of the experimental setup can be found in Appendix A.3.6.
@@
To attack a target node, an adversary node needs to share an edge with the target, either as predecessor or as successor. If the adversary is the successor of the target, then it simply drops packets coming from a target without registering them in its Bloom filter. If the adversary is the predecessor of a target, then it registers the received packet tags in the Bloom filter but then drops packets directed to the target, instead of actually sending them. The purpose of the attack is to degrade the measured reliability of target nodes compared to their true reliability. Thus, the adversary is more successful as the protocol more heavily underestimates the true reliability of target nodes. The cost of the attack for the adversary is an ‘opportunity cost’ given by measured reliability degradation suffered by adversarial nodes due to dropping packets for the attack. The attack is thus more costly as the protocol estimates a lower reliability for adversarial nodes, compared to the reliability that those nodes would have if they did not engage in an attack.
@@
We run over five hundred simulations with various combinations of $|A|$ and $|T|$. In each experiment we compute the adversarial cost for the $|A|$ malicious nodes $c_{A}=|A|-\sum_{i\in A}\hat{\rho}_{i}$ and the cost imposed on the $|T|$ targets $c_{T}=|T|-\sum_{j\in T}\hat{\rho}_{j}$. Note that the overall attack scale increases proportionally to $|A|\times|T|$, as this determines the number of links where packets are dropped and consequently the magnitude of the attack costs $c_{A}$ and $c_{T}$. This results in $c_{A}$ and $c_{T}$ values ranging from below $10^{-2}$ to above $10$. We use log scale to represent results for the full range of values from small to large.
@@
Figure 3: Results for adversarial settings where all honest nodes are reliable. Each sample represents the aggregate reliability penalties $c_{T}$ for $1\leq|T|\leq 64$ targets ($y$ axis) and $c_{A}$ for $1\leq|A|\leq 64$ adversaries ($x$ axis) in each simulation run (total $510$ runs).
@@
In Figure 3 we show the result of each simulation as a dot with coordinates $(c_{A},c_{T})$, As we can see, all the results are close to the diagonal $c_{A}=c_{T}$, meaning that by engaging in this attack a set of malicious nodes take a combined reliability penalty similar to the one they impose on the set of targets, even in scenarios where $40\%$ of the targets’ predecessors and/or successors are adversarial. In the experiments shown in this section we consider that all honest nodes have perfect reliability and that all the packet drops are due to the attack. Appendix A.3.7 shows additional results where the network includes both adversarial and honest but unreliable nodes.
@@ 4.2 Protocol overhead evaluation
In this section we evaluate the overhead introduced by the proposed protocols in terms of the amount of data that needs to be broadcast in each epoch, the data that needs to be stored long-term in a blockchain, and the computational overhead of VRF-based routing. We consider that a number of variables necessary for our protocols are already publicly available for the basic functioning of the mixnet, even before any reliability measurement protocols are in place. This includes: parameters defining the mixnet topology, per-epoch selection and assignment of mix nodes to layers, node public keys (both for signing broadcast information and for decryption of packets), node rewards, client subscriptions and information on spent client credentials. We thus focus on the additional overhead introduced by the proposed reliability estimation protocol, summarized in Table 2.
We consider a setup with $100$ million packets routed per epoch, of which $1$ million ($1\%$) are measurement packets. The mixnet includes $80$ gateways and $240$ mix nodes divided in $3$ layers of $80$ nodes. The data that needs to be broadcast in each epoch for computing reliability scores includes:
@@
We consider that the computed reliability scores for all nodes and gateways are stored long-term in the blockchain to act as a historical record of their measured reliability. Considering a precision of four decimal places, this requires $2$ bytes per node and gateway, resulting in $640$ bytes for $240$ nodes plus $80$ gateways each epoch. Additional measurement protocol parameters that need to be available in the blockchain, such as $p_{\mathbf{lot}}$ and $\bar{\tau}$, may take a few extra bytes but do not need to be updated every epoch.
@@
Node reliability scores 2 B 320 640 B
Table 2: Overhead reliability measurement protocol
@@ 5 Related Work
Mixnets for voting applications. A key use case for mixnets in the literature is voting applications [30, 37], where the mixnet must ensure both the unlinkability of voters to their votes, as well as the integrity of the final tally. Therefore, in addition to privacy properties (input-output unlinkability), these solutions offer very high levels of assurance in terms of verifiability of correct behaviour by all participants, e.g.: each client is authorized by a credential to send a single well-formed ballot and verified to do so, while intermediary mix nodes are guaranteed to correctly process every packet without dropping, adding, or substituting any packets. This level of assurance is however achieved at a significant cost per routed packet: first, all transmissions between any parties happen “in public” via a broadcast channel (or “bulletin board”), so that everyone can verify who sent which (encrypted) packets at each step and no one can deny having received any packets (because they are all public); furthermore, every mix node must also produce and publish NIZK-based shuffle proofs of correctness for its published output packets [24], so that those outputs are considered valid and further processed by the next mix node along the route. The end-to-end latency and cost of this approach are acceptable in voting but is impractical for general-purpose internet access and other applications,
Mixnets for micro-blogging. Atom [25] is a mixnet designed to send short messages (as in micro-blogging) that provides similar reliability properties using lightweight traps. Atom can route one million $32-$byte messages in $28$ minutes, with latency growing with the number of routed messages and with message size. This is in contrast to the setup considered in this paper, where we expect the mixnet to route client messages with a size in (tens or hundreds of) KB within a second, independently of the number of client messages being routed. This is achieved in our design because the proposed reliability estimation protocols are executed a posteriori, after all the messages of an epoch have been delivered, and thus have no impact on the processing time needed to deliver packets compared to the same mixnet not running our reliability estimation protocols. Although Atom’s amount of broadcast data is not explicitly provided, given its design, the amount of required broadcast data grows with the number of client messages, in contrast to our approach, where the broadcast overhead remains constant for a given network size and measurement accuracy.
@@
Mixnet reliability and accountability. Finally, we review prior work that specifically tackles reliability or accountability. Early research includes the proposal by Danezis and Sassaman [10] of having mix nodes send heartbeat traffic to themselves through the mixnet, similar to the loops of cover traffic implemented in Loopix, to detect active attacks by malicious nodes that drop or substitute messages. The idea of using reputation to increase the reliability of mixnets was already put forward in early work by Dingledine et al. [14], who proposed a solution that relies on semi-trusted third-party witnesses to verify whether mix nodes are correctly processing messages and help clients avoid under-performing nodes when choosing packet paths. This work however explicitly excludes active adversaries that try to compromise reliability measurements. Follow-up work by Dingledine and Syverson [17] eliminates the need for trusted witnesses by using test measurements and failure reports linked to reputation, as well as commitments to communal randomness, to build mix cascades that can detect dropped or substituted packets. This solution however suffers from “creeping death” attacks on measurements enabled by adversaries biasing the measurement protocol in their favor, such that over time honest nodes lose more reputation than adversarial nodes. Note that our solution avoids this problem by penalizing adversarial nodes symmetrically with respect to the reputational damage inflicted on their targets. Finally, more recent proposals by Leibowitz et al. [28] or Boyen et al. [3] rely on either honest mix nodes having perfect reliability (any packet loss is considered malicious), or trusted auditing authorities, namely less realistic assumptions in practical scenarios of interest. Their goal is also accountability which is crucially different from ours: to disconnect nodes that drop packets rather than to estimate their reliability in a decentralized manner. We note that the “trip-wire” technique of [3] bears some similarity to our measurement packets — in the sense that trip-wires are indistinguishable packets aimed at detecting failing relays. Nevertheless, trip-wires do not have the distributional characteristics of measurement packets that are generated during our VRF-based routing and enable us to measure the reliability of relays — trip-wires merely facilitate a single bit output. Furthermore, trip-wires fundamentally rely on trusted auditors - while VRF-based routing does not require trusted parties to work.
6 Conclusion
@@
We have proposed a solution to compute reliability estimates for a mixnet’s links and nodes in an efficient and decentralized manner, given client credentials that authorize sending packets through the mixnet. To realize this concept, we put forward a link reliability estimation protocol that uses covert measurement packets to estimate link reliability for legitimate packets across the mixnet, and the associated primitive of VRF-based routing that ensures that legitimate client packets have routes chosen according to the mixnet’s routing policy, with malicious clients being unable to bias their packets. We further demonstrate how to use the link estimates as input to a node reliability estimation protocol that allows mixnet participants to identify the under-performers amongst them.
@@
Besides formally arguing the security properties of our new primitive, we have empirically evaluated via simulations the practical accuracy of the proposed protocols. To make our proposal practically relevant and easy to integrate with the currently deployed Nym mixnet, we consider $3$ layers and $80$ mix nodes per layer. We find that for this network size one million measurements are sufficient to provide accurate estimations for reliability, with the number of required measurements remaining constant regardless of how many legitimate packets are routed. Our novel VRF-based routing primitive is however applicable to any other decryption mixnet given the two required functionalities, namely issuance of epoch-based credentials for authorized usage and a broadcast channel, e.g., a blockchain.
@@
In contrast to verifiable solutions, our scheme is compatible with low-latency continuous time mix networks such as Nym as the verification steps happen a posteriori and thus do not add latency to the delivery of packets. Furthermore, the broadcast channel is used for a small fraction of packets ($1\%$ or less if the network routes several hundred million packets or more) rather than all of them. For each measurement packet only a few parameter values need to be broadcast to enable verifying the packet header, rather than requiring (as in e-voting oriented solutions) storing the entire packet including its payload, recorded repeatedly at each step of its route. Our approach provides the first scalable low-latency estimates of reliability for real-world general-purpose mixnets composed of decentralized mix nodes, a previously unsolved problem [17].
Appendix A Appendix
A.1 Proof of Theorem 1
@@
We show how to incorporate a challenge for the Decisional Diffie Hellman assumption $\langle g,a,b,c\rangle$ into the encoding calculation. We set the credential key $\alpha$ to be $a$ and the node public-key $y_{i}$ to be equal to $b$. It follows that the value $\tilde{r}_{i}$ can be calculated as a function of the DDH challenge by $H(\mathsf{rnd},c^{e})$ where $e=b_{0}\ldots b_{i-1}\cdot r_{\mathrm{pkt}}$. Conditional on $e\neq 0$, it follows that $\tilde{r}_{i}$ is uniformly random, unless the adversary queries the value $c^{e}$ to the random oracle. Note that this event can only happen with non-negligible probability in the case the challenge $\langle g,a,b,c\rangle$ is a DDH tuple (as in the other case, the value $c$ is random and hence can only be predicted with negligible probability by the adversary). As a result, if the adversary queries $c^{e}$ we can readily build a distinguisher by examining the random oracle queries of the adversary. Specifically, the distinguisher can test each query $q$ for $q=c^{e}$ and output $1$ if this is the case while producing 00 otherwise. It is easy to see that this algorithm will produce $1$ with non-negligible probability when the tuple $\langle g,a,b,c\rangle$ follows the DDH distribution while it will produce 00 with overwhelming probability whenever $\langle g,a,b,c\rangle$ is a random tuple. Finally, observe that the probability that $e=0$ is negligible given the randomness of the constituent values in the random oracle model.
The above argument generalizes via a hybrid argument in a straightforward manner to the case that there are multiple clients sending multiple packets. Note that it is in this step that the VRF property is needed since we want to ensure that the $\alpha_{0}$ values of different packets appear independently sampled from the base group. The pseudorandomness property provides us exactly that: the value $r_{\mathsf{pkt}}$ would randomize $\alpha$ within the base group enabling us to repeat the DDH distinguishing argument for each packet.
@@
(II) The statement asks for the pseudorandomness of the values $r_{0},\ldots,r_{\nu-1}$, even in the case that the values $vk,\alpha,y_{0}$ are adversarially chosen. Note that in this case we cannot rely on the DDH assumption or the properties of the VRF. Instead the key observation is that each value $r_{i}$ depends on the public nonce value via the random oracle as well as some other value $s_{i}$ that as long as it is unique, $r_{i}$ would be an independently sampled value. To prove this we consider the event that there are two values $s=g^{e},s^{\prime}=g^{e^{\prime}}$ in this sequence that are equal, i.e., $e=e^{\prime}$. Depending on the step when these values have been calculated the exponents will be a product of a number of values some of which are committed by the adversary, e.g., the discrete logarithm of the key of an adversarial gateway, the discrete logarithm of a credential’s randomization key $\alpha$ as well as values produced by random oracle evaluations of such values e.g., the blinding exponents $b_{0},b_{1},b_{2},\ldots$ and the keys of the mix-nodes. Conditional on no collisions in the random oracle $H(\cdot)$ or the VRF function, all these individual values involved in $e,e^{\prime}$ are distinct and, in case they are not outputs of the random oracle, are committed prior to the production of $\mathsf{nonce}$. Based on this it follows that the event $e=e^{\prime}$ would imply the event that $z_{1}\cdot z_{k}\cdot h_{1}\ldots h_{t}=1$ where $h_{1},\ldots,h_{t}$ are random oracle outputs that depend on $\mathsf{nonce}$ and $z_{1},\ldots,z_{k}$ values selected by the adversary. It is easy to see via generalized birthday problem [36] that this would require subexponential overhead in the worst-case (note in our setting $t$ is a small constant).
(III) This statement follows easily due to the fact that the only difference between measurement and non-measurement packets is in the calculation of the $\alpha_{0}$ value: in the case of a non-measurement packet this value is equal to $\alpha^{r_{\mathsf{pkt}}}$, while in the case of a measurement packet it is equal to $g^{r_{\mathsf{pkt}}}$. In the view of a relay node, the exponentiation by the (hidden) value $r_{\mathsf{pkt}}$ results in a full randomization of the resulting element within the base group (with the precondition $\alpha\neq 1$); as a result the two cases areco indistinguishable assuming the security of the VRF function. ∎
A.2 Approaches to setting $\hat{\beta}_{e}$